# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ 5. $(\mathrm{5})$
We first prove the first equivalence using the definition of $\mathcal{S}(\Pi)$.
$\zeta\in\mathcal{S}(\Pi)\Rightarrow\zeta=\Pi\zeta\Pi=\Pi\Pi\zeta\Pi=\Pi\zeta.$ (50)
$\zeta=\Pi\zeta\Rightarrow\zeta=\zeta\Pi\Rightarrow\zeta=\Pi\zeta\Pi\Rightarrow\zeta\in\mathcal{S}(\Pi).$ (51)
This completes the proof of first equivalence. Let $k_{\zeta,i}\neq 0$ denote the eigenvalue corresponding to eigenvector $|\phi_{i}\rangle$, then $\forall i$
$\Pi\zeta=\zeta\Rightarrow\Pi\zeta|\phi_{i}\rangle=\zeta|\phi_{i}\rangle\Rightarrow k_{\zeta,i}\Pi|\phi_{i}\rangle=k_{\zeta,i}|\phi_{i}\rangle\Rightarrow\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall\ i.$ (52)
Now taking from the right side,
$\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall i$ $\Rightarrow\sum_{i}k_{\zeta,i}\Pi|\phi_{i}\rangle\langle\phi_{i}|\Pi=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ (53)
$\Rightarrow\Pi\left(\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|\right)\Pi=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ (54)
$\Rightarrow\Pi\zeta\Pi=\zeta$ (55)
$\Rightarrow\zeta\in\mathcal{S}(\Pi).$ (56)
Thus $\zeta\in\mathcal{S}(\Pi)\Leftrightarrow\Pi\zeta=\zeta\Leftrightarrow\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall i$. ∎
Lemma A.5.
Given $\nu\in\mathcal{P}(\mathcal{H}),$$\langle\phi|\nu|\phi\rangle\geq\|\nu\|_{\infty,0}\ \forall\ |\phi\rangle:\Pi_{\nu}|\phi\rangle=|\phi\rangle$ and $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\text{ at }\Pi_{\nu}|\phi\rangle=|\phi\rangle\Leftrightarrow\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$.
Proof.
We first prove the inequality and then show the equivalence of equality by showing the following 3 conditions as equivalent.
  1. 1.
    $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty,0}$,
  2. 2.
    $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}$,
  3. 3.
    $\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$.
Using eigenvalue composition of $\nu$,
$\langle\phi|\nu|\phi\rangle=\sum_{i}\langle\phi|k_{\nu,i}|e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\langle\phi||e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}.$ (57)
Note that $k_{\nu,i}\geq\|\nu\|_{\infty,0}\ \forall i$ and
$\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\sum_{i}\langle\phi|e_{i}\rangle\langle e_{i}|\phi\rangle=\langle\phi|\Pi_{\nu}|\phi\rangle=\langle\phi|\Pi_{\nu}\Pi_{\nu}^{\min}|\phi\rangle=\langle\phi|\Pi_{\nu}^{\min}|\phi\rangle=\langle\phi\|\phi\rangle=1.$
Hence,
$\langle\phi|\nu|\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}\geq\sum_{i}\|\nu\|_{\infty,0}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\|\nu\|_{\infty,0}\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\|\nu\|_{\infty,0}.$ (58)
This proves the inequality of lemma.
Now equality holds iff $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}$. This proves the equivalence with statement 2. Note that $\Pi_{\nu}|\phi\rangle=|\phi\rangle$ and so $|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle$. On substituting $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}$, we get
$|\phi\rangle=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty,0}}|e_{i}\rangle\langle e_{i}|\phi\rangle=\Pi_{\nu}^{\min}|\phi\rangle.$ (59)
For showing the equivalence, converse is also needed, which is proved below.
$|\phi\rangle=\Pi_{\nu}^{\min}|\phi\rangle=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty,0}}|e_{i}\rangle\langle e_{i}|\phi\rangle\Rightarrow\langle e_{i}|\phi\rangle=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}.$ (60)
So, we conclude $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty,0}\text{ at }\Pi_{\nu}|\phi\rangle=|\phi\rangle\Leftrightarrow\langle e_{i}|\phi\rangle=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}\Leftrightarrow\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$. Hence, $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty,0}\Leftrightarrow\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$. ∎
Lemma A.6.
For any $\nu\in\mathcal{P}(\mathcal{H})$, $\mathrm{Tr}\left(\zeta\nu\right)\geq\|\nu\|_{\infty,0}\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu})$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\Leftrightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$.
Proof.
The proof is done in 2 steps. First we have shown that $\mathrm{Tr}\left(\zeta\nu\right)\geq\|\nu\|_{\infty,0}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$, followed by showing $\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}$. Combining these two, we get the desired result.
  1. 1.
    Take eigenvalue decomposition of $\zeta=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ with $\sum_{i}k_{\zeta,i}=1$. Then $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle$. From Lemma A.4, we know that $\Pi_{\nu}|\phi_{i}\rangle=|\phi_{i}\rangle$. Now, using Lemma A.5, $\langle\phi_{i}|\nu|\phi_{i}\rangle\geq\|\nu\|_{\infty,0}$ and $\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty,0}\Leftrightarrow\Pi_{\nu}^{\min}|\phi_{i}\rangle=|\phi_{i}\rangle$. Hence,
    $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle\geq\|\nu\|_{\infty,0}\sum_{i}k_{\zeta,i}=\|\nu\|_{\infty},$ (61)
    with $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\Leftrightarrow\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty,0}\ \forall\ i\Leftrightarrow\Pi_{\nu}^{\min}|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall\ i$. Hence,
    $\Pi_{\nu}^{\min}\zeta\Pi_{\nu}^{\min}=\sum_{i}k_{\zeta,i}\Pi_{\nu}^{\min}|\phi_{i}\rangle\langle\phi_{i}|\Pi_{\nu}^{\min}=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|=\zeta\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min}).$ (62)
    So, $\mathrm{Tr}\left(\zeta\nu\right)\geq\|\nu\|_{\infty,0}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$.
  2. 2.
    Using statement $(\mathrm{8})$ of Lemma A.1, we get
    $\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}.$ (63)
Combining the two, we get the desired result. ∎
Appendix B Proof of Theorem III.1
@@
$e(\Lambda)$ can be written as $e(\Lambda)=\left(1+R_{\Lambda,p}(\rho,\sigma)\right)^{-1}$, where
@@
On taking $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\Lambda_{\sigma}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$,
@@
where $(a)$ is due to $\Pi_{\rho}=\Pi_{\sigma}$. Also, $\mathrm{Tr}\left(\Lambda_{\rho}^{s}\right)=\mathrm{Tr}(\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2})=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right).$ Similarly, $\mathrm{Tr}\left(\sigma^{-1/2}\Lambda_{\sigma}^{s}\sigma^{-1/2}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ and $\mathrm{Tr}\left(\Lambda_{\sigma}^{s}\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$. On substituting these in $R_{\Lambda}(p_{\rho}\rho,p_{\sigma}\sigma)$, we get
@@
Using Lemma A.3 and Lemma A.6 from Appendix A, we get
@@
$\mathrm{Tr}\left(\Lambda_{\sigma}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)\geq\|\sigma^{-1/2}\rho\sigma^{-1/2}\|_{\infty,0}\text{ with equality iff }\Lambda_{\sigma}^{s,c}\in\mathcal{S}(\mathrm{T}^{\min}).$
Substituting these bounds in the expression of $R_{\Lambda,p}(\rho,\sigma)$ in Equation (65), we get upper bound as
@@
=$=\begin{cases}c_{\rho}>0,c_{\sigma}=0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ c_{\rho}=0,c_{\sigma}>0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ c_{\rho}\geq 0,c_{\sigma}\geq 0,(c_{\rho},c_{\sigma})\neq(0,0),&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}={R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}.\end{cases}$ (68)
@@
Further, the minimum of $e(\Lambda)$ being achieved at $c_{\rho}$ and $c_{\sigma}$ satisfying (68) and $\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})$ and $\Lambda_{\sigma}^{s,c}\in\mathcal{S}(\mathrm{T}^{\min})$. Combining $\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})$ with $c_{r}=0$ and $c_{r}>0$ gives $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0$ and $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ respectively. Similarly, we get condition on $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$. Thus, the final condition for equality in (69) and so minimum of $e(\Lambda)$ is obtained as
@@
which is the desired result.
Appendix C Generalized form of operator given the projector
Lemma C.1.
For any $\zeta\in\mathcal{P}(\mathcal{H})$ and projector $\Pi$, the following three are equivalent:
$(\mathrm{1})\ \Pi\zeta\Pi=0\qquad(\mathrm{2})\ \Pi\zeta=\zeta\Pi=0\qquad(\mathrm{3})\ \zeta\in\mathcal{P}(\mathrm{I}-\Pi).$
@@ Appendix C Generalized form of operator given the projector
We will prove $(\mathrm{1})\Rightarrow(\mathrm{2})\Rightarrow(\mathrm{3})\Rightarrow(\mathrm{1})$.
(1$\to$(2): Let $\zeta=\sum_{i}k_{i}|\psi_{i}\rangle\langle\psi_{i}|$ be eigenvalue decomposition of $\zeta$ with $k_{i}>0$, then
$\Pi\zeta\Pi=0\Rightarrow\mathrm{Tr}\left(\Pi\zeta\right)=0\Rightarrow\sum_{i}k_{i}\langle\psi_{i}|\Pi|\psi_{i}\rangle=0.$
Note that $k_{i}>0$ and $\langle\psi_{i}|\Pi|\psi_{i}\rangle\geq 0$, hence $\sum_{i}k_{i}\langle\psi_{i}|\Pi|\psi_{i}\rangle=0$ iff $\langle\psi_{i}|\Pi|\psi_{i}\rangle=0$. So we obtain
$\Pi\zeta\Pi=0\Rightarrow\langle\psi_{i}|\Pi|\psi_{i}\rangle=0\ \forall\ i\stackrel{{\scriptstyle(a)}}{{\Rightarrow}}\Pi|\psi_{i}\rangle=0\ \forall\ i\ \Rightarrow\Pi\zeta=\zeta\Pi=0.$
$(a)$ follows from the fact that, if norm squared of a vector $\Pi|\psi_{i}\rangle$, which is $\langle\psi_{i}|\Pi|\psi_{i}\rangle=0$, then the vector $\Pi|\psi_{i}\rangle=0$.
(2$\to$(3): $\Pi\zeta=\zeta\Pi=0\Rightarrow(\mathrm{I}-\Pi)\zeta=\zeta=\zeta(\mathrm{I}-\Pi)\Rightarrow\zeta=(\mathrm{I}-\Pi)\zeta(\mathrm{I}-\Pi)\Rightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi)$.
(3$\to$(1): $\zeta\in\mathcal{P}(\mathrm{I}-\Pi)\Rightarrow\zeta=(\mathrm{I}-\Pi)\zeta(\mathrm{I}-\Pi)\Rightarrow\Pi\zeta\Pi=\Pi(\mathrm{I}-\Pi)\zeta(\mathrm{I}-\Pi)\Pi=(\Pi-\Pi)\zeta(\Pi-\Pi)=0$. So, we get equivalence of three statements. ∎
Lemma C.2.
For any $\zeta\in\mathcal{P}(\mathcal{H})$ and $\nu\in\mathcal{P}(\mathcal{H})$,
$\mathrm{Tr}\left(\zeta\nu\right)=0\Leftrightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu}).$
Proof.
Note that $\zeta,\nu\in\mathcal{P}(\mathcal{H})$, hence we have $\nu^{1/2}\zeta\nu^{1/2}\geq 0$. Now
$\mathrm{Tr}\left(\zeta\nu\right)=0\stackrel{{\scriptstyle(a)}}{{\Leftrightarrow}}\mathrm{Tr}\left(\nu^{1/2}\zeta\nu^{1/2}\right)=0\stackrel{{\scriptstyle(b)}}{{\Leftrightarrow}}\nu^{1/2}\zeta\nu^{1/2}=0\stackrel{{\scriptstyle(c)}}{{\Leftrightarrow}}\Pi_{\nu}\zeta\Pi_{\nu}=0\stackrel{{\scriptstyle(d)}}{{\Leftrightarrow}}\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu}).$
$(a)$ follows from cyclic property of $\mathrm{Tr}\left(\cdot\right)$ operator. $(b)$ follows from the fact that $\nu^{1/2}\zeta\nu^{1/2}\geq 0$. $(c)$ follows from multiplying both side by $\nu^{-1/2}$ or $\nu^{1/2}$ (for reverse direction). $(d)$ utilizes Lemma C.1. ∎
Lemma C.3.
@@
By taking hermitian and multiplying with $\nu^{-1}$ and $\nu$, we obtain
@@
Now, beginning with $\zeta\in\mathcal{P}(\Pi_{\nu})$, we get
@@
Using the definition of projector, we obtained that $\mathrm{P}=\Pi_{\nu^{-1}\Pi\nu^{-1}}=\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}\left(\nu^{-1}\Pi\nu^{-1}\right)$ and $\nu^{-1}\Pi_{\nu}=\nu^{-1}$. Now beginning wit $\mathrm{P}\Pi_{\nu}$, we get
$\mathrm{P}\Pi_{\nu}$ $=\Pi_{\nu^{-1}\Pi\nu^{-1}}\Pi_{\nu}$
$=(\nu^{-1}\Pi\nu^{-1})^{-1}(\nu^{-1}\Pi\nu^{-1})\Pi_{\nu}$
$=(\nu^{-1}\Pi\nu^{-1})^{-1}\nu^{-1}\Pi\nu^{-1}\Pi_{\nu}$
$=(\nu^{-1}\Pi\nu^{-1})^{-1}\nu^{-1}\Pi\nu^{-1}=\mathrm{P},$
@@
So we have $\Pi\nu\mathrm{P}\nu=\nu\mathrm{P}\nu$. Now, using the Lemma C.3, we obtain $\zeta\in\mathcal{P}(\Pi_{\nu\mathrm{P}\nu})\Rightarrow\zeta\in\mathcal{P}(\Pi).$
Lemma C.6.
@@
Note that $\zeta\in\mathcal{P}(\mathrm{P})$ and $\mathrm{P}\Pi_{\nu}=\mathrm{P}$, hence
@@
Now, to prove $\nu^{-1}\zeta\nu^{-1}\in\mathcal{P}(\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}})$,
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}(\nu^{-1}\mathrm{P}\nu^{-1})\nu\zeta\nu(\nu^{-1}\mathrm{P}\nu^{-1})=\nu^{-1}\zeta\nu^{-1}.$
$(a)$ is obtained from the fact that $\eta\Pi_{\eta}=\Pi_{\eta}\eta=\eta$ on taking $\eta=\nu^{-1}\mathrm{P}\nu^{-1}$. Thus we get $\nu^{-1}\zeta\nu^{-1}\in\mathcal{P}(\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}})$, which completes the proof. ∎
Theorem C.1.
Given $\Pi,\nu\in\mathcal{P}(\mathcal{H}),\Pi\Pi_{\nu}=\Pi$, then the following statements are equivalent

### v2

@@ Appendix A Proof of Theorem 1
The proof follows similar to the proof of Lemma A.1. ∎
@@
Proof of Theorem 1: From the definition of $e(\Lambda)$ in (4), $e(\Lambda)$ can be written as $e(\Lambda)=\left(1+R_{\Lambda,p}(\rho,\sigma)\right)^{-1}$, where
@@
On taking $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\Lambda_{\sigma}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$, we get
@@
where $(a)$ is due to $\Pi_{\rho}=\Pi_{\sigma}$. Also, $\mathrm{Tr}\left(\Lambda_{\rho}^{s}\right)=\mathrm{Tr}(\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2})=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right).$ Similarly, $\mathrm{Tr}\left(\sigma^{-1/2}\Lambda_{\sigma}^{s}\sigma^{-1/2}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ and $\mathrm{Tr}\left(\Lambda_{\sigma}^{s}\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$. On substituting these in $R_{\Lambda,p}(\rho,\sigma)$, we get
@@
Using Lemma A.1 and Lemma A.2, we get
@@
$\mathrm{Tr}\left(\Lambda_{\sigma}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)\geq\|\sigma^{-1/2}\rho\sigma^{-1/2}\|_{\infty,0}\text{ with equality iff }\Lambda_{\sigma}^{s,c}\in\mathcal{S}(\mathrm{T}^{\min}),$
where $\mathrm{T}^{\max}=\Pi_{\sigma^{-1/2}\rho\sigma^{-1/2}}^{\max}$ and $\mathrm{T}^{\min}=\Pi_{\sigma^{-1/2}\rho\sigma^{-1/2}}^{\min}$. Substituting these bounds in the expression of $R_{\Lambda,p}(\rho,\sigma)$ in (43), we get
@@
$=\begin{cases}c_{\rho}>0,c_{\sigma}=0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ c_{\rho}=0,c_{\sigma}>0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ c_{\rho}\geq 0,c_{\sigma}\geq 0,(c_{\rho},c_{\sigma})\neq(0,0),&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}={R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}.\end{cases}$ (46)
@@
Further, the minimum of $e(\Lambda)$ being achieved at $c_{\rho}$ and $c_{\sigma}$ satisfying (46) and $\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})$ and $\Lambda_{\sigma}^{s,c}\in\mathcal{S}(\mathrm{T}^{\min})$. Combining $\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})$ with $c_{\rho}=0$ and $c_{\rho}>0$ gives $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0$ and $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ respectively. Similarly, we get condition on $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$. Thus, the final condition for equality in (47), and so the condition on $\Lambda$ for the minimum of $e(\Lambda)$ to be achieved, is obtained as
@@
which completes the proof.
Appendix B Generalized form of operator $\zeta$ given a constraint on subspace where it lies
Lemma B.1.
For a variable operator $\zeta\in\mathcal{P}(\mathcal{H})$ and a fixed operator $\nu\in\mathcal{P}(\mathcal{H})$, the following are equivalent:
$(\mathrm{1})\ \mathrm{Tr}\left(\zeta\nu\right)=0\qquad(\mathrm{2})\ \Pi_{\nu}\zeta=\zeta\Pi_{\nu}=0\qquad(\mathrm{3})\ \zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu})\qquad(\mathrm{4})\ \Pi_{\nu}\zeta\Pi_{\nu}=0.$
@@ Appendix B Generalized form of operator $\zeta$ given a constraint on subspace where it lies
We will prove $(\mathrm{1})\Rightarrow(\mathrm{2})\Rightarrow(\mathrm{3})\Rightarrow(\mathrm{4})\Rightarrow(\mathrm{1})$.
[…]
$\mathrm{Tr}\left(\zeta\nu\right)=0\Rightarrow\sum_{i}k_{i}\langle\psi_{i}|\nu|\psi_{i}\rangle=0.$
Note that $k_{i}>0$ and $\langle\psi_{i}|\nu|\psi_{i}\rangle\geq 0$, hence $\sum_{i}k_{i}\langle\psi_{i}|\nu|\psi_{i}\rangle=0$ iff $\langle\psi_{i}|\nu|\psi_{i}\rangle=0$. So we obtain
$\langle\psi_{i}|\nu|\psi_{i}\rangle=0\ \forall\ i\stackrel{{\scriptstyle(a)}}{{\Rightarrow}}\nu^{1/2}|\psi_{i}\rangle=0\ \forall\ i\ \Rightarrow\Pi_{\nu}|\psi_{i}\rangle=0\ \forall\ i\ \Rightarrow\Pi_{\nu}\zeta=\zeta\Pi_{\nu}=0.$
$(a)$ follows from the fact that, if $\|\nu^{1/2}|\psi_{i}\rangle\|^{2}=\langle\psi_{i}|\nu|\psi_{i}\rangle=0$, then the vector $\nu^{1/2}|\psi_{i}\rangle=0$.
[…]
Lemma B.2.
@@
By taking Hermitian on both sides and multiplying both sides with $\nu^{-1}$ and $\nu$, we obtain
@@
Now, for any $\zeta\in\mathcal{P}(\Pi_{\nu})$, we get
@@
Using the definition of projector, we obtained that $\mathrm{P}=\Pi_{\nu^{-1}\Pi\nu^{-1}}=\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}\left(\nu^{-1}\Pi\nu^{-1}\right)$ and $\nu^{-1}\Pi_{\nu}=\nu^{-1}$. Now,
$\mathrm{P}\Pi_{\nu}$ $=\Pi_{\nu^{-1}\Pi\nu^{-1}}\Pi_{\nu}=(\nu^{-1}\Pi\nu^{-1})^{-1}(\nu^{-1}\Pi\nu^{-1})\Pi_{\nu}$
$=(\nu^{-1}\Pi\nu^{-1})^{-1}\nu^{-1}\Pi\nu^{-1}\Pi_{\nu}=(\nu^{-1}\Pi\nu^{-1})^{-1}\nu^{-1}\Pi\nu^{-1}=\mathrm{P},$
@@
Therefore, we have $\Pi\nu\mathrm{P}\nu=\nu\mathrm{P}\nu$. Now, using the Lemma B.2, we obtain $\zeta\in\mathcal{P}(\Pi_{\nu\mathrm{P}\nu})\Rightarrow\zeta\in\mathcal{P}(\Pi).$
Lemma B.5.
@@
Note that $\zeta\in\mathcal{P}(\mathrm{P})$ and $\mathrm{P}\Pi_{\nu}=\mathrm{P}$, hence,
@@
Now, to prove $\nu^{-1}\zeta\nu^{-1}\in\mathcal{P}(\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}})$, we show that
@@
$\stackrel{{\scriptstyle(a)}}{{=}}(\nu^{-1}\mathrm{P}\nu^{-1})\nu\zeta\nu(\nu^{-1}\mathrm{P}\nu^{-1})=\nu^{-1}\zeta\nu^{-1}.$
$(a)$ is obtained from the fact that $\eta\Pi_{\eta}=\Pi_{\eta}\eta=\eta$ on taking $\eta=\nu^{-1}\mathrm{P}\nu^{-1}$. Thus, we get $\nu^{-1}\zeta\nu^{-1}\in\mathcal{P}(\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}})$, which completes the proof. ∎
Theorem B.1.
Given $\Pi,\nu\in\mathcal{P}(\mathcal{H}),\Pi\Pi_{\nu}=\Pi$, then the following statements regarding constraints on $\zeta\in\mathcal{P}(\mathcal{H})$ are equivalent